OpenStudy (anonymous):

If x^3+y^3=9 and dx/dt=4, find dy/dt when x=2

OpenStudy (sierra!):

I think its -16

OpenStudy (anonymous):

how did you the answer? can you please explain

OpenStudy (campbell_st):

that is the correct answer

OpenStudy (sierra!):

when x = 2, y^3 = 9 - 8 = 1==> y = 1 differentiating implicitly 3x^2 dx/dt + 3y^2dy/dt = 0 =>x^2 dx/dt + y^2 dy/dt = 0 substitute x = 2, y = 1 and dx/dt = 4 (4)(4) + (1)dy/dt = 0 dy/dt = -16

OpenStudy (sierra!):

Can i have a medal

OpenStudy (anonymous):

thank you!

OpenStudy (sierra!):

No thank you